OpenStudy (anonymous):

If s(x) = 2 – x2 and t(x) = 3x, which value is equivalent to s(t(-7)) –439 –141 153 443

OpenStudy (solomonzelman):

first you need to find the s(t(x)).

OpenStudy (solomonzelman):

plug in 3x instead of x, into 2-x²

OpenStudy (solomonzelman):

u get?

OpenStudy (anonymous):

-7

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (solomonzelman):

s(x) = 2 – x² correct?

OpenStudy (anonymous):

correct

OpenStudy (solomonzelman):

ok, if I wanted to find s(A) wat would I do?

OpenStudy (anonymous):

s(A)? I dont know

OpenStudy (solomonzelman):

s(x)=2-x² s(A)=2-(A)² just sub A for x

OpenStudy (solomonzelman):

what if I wanted to find s(3), what would I do?

OpenStudy (anonymous):

sub x for 3?

OpenStudy (solomonzelman):

yes sub 3 for x

OpenStudy (solomonzelman):

s(x)=2-x² s(3) = ?

OpenStudy (anonymous):

-7?

OpenStudy (anonymous):

right?

OpenStudy (solomonzelman):

yes

OpenStudy (solomonzelman):

so s(x)=2-x² what is s(3x) ?

OpenStudy (anonymous):

sorry back was helping my dad

OpenStudy (anonymous):

2-3x^2?

OpenStudy (solomonzelman):

i am helping another student too, so it's ok

OpenStudy (solomonzelman):

s(3x)=2-(3x)² like this

OpenStudy (anonymous):

okay haha

OpenStudy (solomonzelman):

s(x-3)=?

OpenStudy (anonymous):

x

OpenStudy (anonymous):

ugghh im confused

OpenStudy (anonymous):

@SolomonZelman